OpenStudy (anonymous):

WILL FAN AND MEDAL The wavelength of a specific photon in the infrared spectrum is 4 × 10^-6 meters. Given that the speed of light is 2.998 × 10^8 m/s, what would be the frequency of this infrared wave? 11.992*10^2 Hz 0.7495*10^-14 Hz 11.992 * 10^-2 Hz 0.7495 * 10^14 Hz

OpenStudy (cuanchi):

speed of light = wavelength x frequency frequency = speed of light/ wavelength do the math!!

OpenStudy (anonymous):

@Cuanchi if I knew how to do the math I wouldn't be on here

OpenStudy (photon336):

@whateven we need to understand the units first, I agree with what @Cuanchi said

OpenStudy (photon336):

speed of light is m/s and the wavelength is in meters.

OpenStudy (anonymous):

This is what my question is I can't change it and if I don't answer correctly my gpa goes down

OpenStudy (photon336):

let's take @Cuanchi's formula \[\frac{ speed of light }{ wavelength } = \frac{ m/s }{ m } = \frac{ 1}{ s } = frequency\] all you need to do is plug in the numbers.

OpenStudy (anonymous):

I don't know where to plug them in and get a result that matches my choices that's the problem

OpenStudy (cuanchi):

The units are wavelength = meters speed of light = meters/ seconds frequency = 1/seconds you dont need to do any conversion of units, just apply the formula that correlate the speed of light with the wavelength and the frequency

OpenStudy (photon336):

\[\frac{ 1 }{ s } = hertz \]

OpenStudy (cuanchi):

it is a simple division of the value of speed of light divided by the wavelenght

OpenStudy (cuanchi):

Hz = 1/seconds = 1/s

OpenStudy (photon336):

plug the numbers in and tell us what you get before you submit your answer

OpenStudy (anonymous):

This still isn't helping. I plug it in like you say and my calculator says syntax error ?

OpenStudy (cuanchi):

2.998 × 10^8 m/s /4 × 10^-6 meters= x.xxx 10^xx

OpenStudy (anonymous):

that's what i typed in. It said syntax error

OpenStudy (photon336):

@whateven use ( ) around the numbers like this \[\frac{ (2.99x10^{8}) m/s }{ (4.0x10^{-6})m }= Hz \]

OpenStudy (photon336):

You can also enter this in on a scientific calculator online

OpenStudy (anonymous):

7.495*10^13 so something is wrong @photon336

OpenStudy (anonymous):

????

OpenStudy (photon336):

That's correct,

OpenStudy (photon336):

but in the answers they wrote it strangely, if you moved the decimal place in the 7.4 1 over to the right what answer choice would you get?

OpenStudy (anonymous):

The last one?

OpenStudy (photon336):

yeah that's it

OpenStudy (anonymous):

Thanks!